# PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection

### v5

@@ Abstract
Anomaly detection in time series analysis is a pivotal task, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios. While prior studies have largely employed reconstruction-based approaches, which limits the models’ representational capacities. Moreover, existing deep learning-based methods are not sufficiently lightweight. Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection. With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 3.2MB. Its efficacy is demonstrated by state-of-the-art results across 9 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms. PatchAD significantly improves the classical F1 score by 50.5%, the Aff-F1 score by 7.8%, and the AUC by 10.0%. The code is publicly available.  [^1] https://github.com/EmorZz1G/PatchAD
1 Introduction
@@
Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from expected behaviour [1, 2, 3]. With the rapid progression of large-scale sensing technologies and storage capabilities, temporal anomaly detection has found widespread application in various sectors. For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [4, 5]. Unsupervised algorithms, capable of operating without labels, have gained wide popularity. These include reconstruction-based[6, 7], density-based[8, 9], contrastive learning[10, 11], autoregression-based[12, 13, 14], large language models (LLMs)[15] etc. However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms. Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature. Effective anomaly detection requires models that can learn the interrelationships between various data channels. Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features. Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data. Specifically, several algorithms have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning. Additionally, they either do not explicitly model inter-channel dependencies [10] or rely solely on basic MLPs or Transformers that focus on temporal information [16], mainly considering point-level features [17].
@@
On the other hand, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points. This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not. This fundamental principle forms the basis of our proposed anomaly detection method.
@@
In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection. Specifically, proposed Multi-scale Patching and Embedding decomposes the input data along the temporal dimension into non-overlapping patches, extracting semantically richer features beyond simple point-level characteristics. Besides, various submodels address various time scales, each focusing on different patch sizes. Patch Mixer Encoders are introduced to learn the complex relationships within and between patches and channels. A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities. In detail, this work makes the following contributions:
@@
Figure 2: Model sizes.
(1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method. PatchAD achieves superior performance on 9 benchmark datasets, surpassing over 30 baselines across various evaluation metrics, with and without prior association.
(2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models.
(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by a theoretical analysis that elucidates the model’s underlying principles and design choices.
2 Preliminaries
2.1 Problem Formulation
@@
2.2 Time Series Anomaly Detection
@@
The time series anomaly detection methods encompass statistical approaches, classical machine learning methods, and deep learning techniques[18]. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [19] and Support Vector Data Description (SVDD) [20]. Deep learning methods include DAGMM, OmniAnomaly [7], USAD [17], LSTM-VAE [6], and DeepAnT [21], which are trained on normal data and detect anomalies by inference using reconstruction errors. Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [22, 23]. While [24] designed six different anomaly injection methods to train the model in a self-supervised approach, [25] utilized outlier exposure and mixup to construct a self-supervised task. Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [26] introduced sequence variational autoencoder model based on smoothness constraints, whereas [27] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance. Recently approaches based on diffusion models and Large language model have also been proposed [28, 15].
2.3 MLP Mixer & Contrastive Learning for Time Series Analysis
@@
The MLP Mixer initially proposed as a novel architecture for computer vision [29], has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like [30, 31, 32] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF).
@@
Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs. Additionally, methods like BYOL [33] and SimSiam [34] have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as [11] and [35] have explored integrating contrastive learning into time series analysis. [10], on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
3 Proposed Method
3.1 Overall Architecture
@@
3.2 Multi-Scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. This enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ \mathcal{X}^{C\times T}\rightarrow\mathcal{X}^{C\times N\times P},$ (1)
where $\operatorname{PE}(\cdot)$ is positional embedding. PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. And $\operatorname{VE}(\cdot)$ is a single linear layer.
$\mathcal{X}_{inter}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times N\times D},$ (2)
$\mathcal{X}_{intra}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times P\times D}.$
3.3 Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. The MLP Mixer consists of two fully connected layers, a normalization layer and a GELU non-linear activation layer, and utilizes a residual connection from input to output. This can be represented as:
$\mathcal{X}=\operatorname{FC}(\operatorname{GELU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (3)
The following are the different types of MLP Mixers utilized in PatchAD:
@@ 3.3 Patch Mixer Encoder
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation. In the end, two separate simple MLPs are utilized to reconstruct the original data, which are then combined.
$\hat{\mathcal{X}}_{1}=\operatorname{Rec1}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec2}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (4)
3.4 Dual Project Constraint

### v6

@@ Abstract
Time series anomaly detection is a pivotal task in data analysis, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios. While prior studies have largely employed reconstruction-based approaches, which limit the models’ representational capacities. Moreover, existing deep learning-based methods are not sufficiently lightweight. Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection. With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 0.403M parameters. Its efficacy is demonstrated by state-of-the-art results across 8 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms. PatchAD significantly improves the classical F1 score by 6.84%, the Aff-F1 score by 4.27%, and the V-ROC by 2.49%. Simultaneously, an in-depth analysis of the mechanisms underlying PatchAD has been conducted from both theoretical and experimental perspectives, validating the design motivations of the model. The code is publicly available at https://github.com/EmorZz1G/PatchAD.
Index Terms:
Time series, anomaly detection, data mining, lightweight, MLP.
I Introduction
@@
Time series anomaly detection (TSAD) is a pivotal aspect of data mining, focused on identifying abnormal patterns within time series data that significantly deviate from expected behaviour [1, 2, 3, 4]. In the era of big data, with the rapid advancement of large-scale sensing technologies and the continuous enhancement of storage capabilities, TSAD has been widely applied in diverse sectors. For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [5, 6]. Unsupervised algorithms, capable of operating without labels, have gained wide popularity [7]. These include reconstruction-based [8, 9], density-based [10], contrastive learning [11, 12], autoregression-based [13, 14, 15], large language models (LLMs) [16] etc. However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms. Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature. Effective anomaly detection requires models that can learn the inter-relationships between various data channels. Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features. Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data. Specifically, several algorithms [17, 11, 18, 19] have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning. For instance, Xu et al. [17] devised a contrastive learning approach that leverages prior correlation and series correlation mining to elucidate the relationship between normal and anomalous patterns. However, this necessitated the employment of a maximization-minimization optimization strategy to train the model and prevent its degradation. Similarly, Yang et al. [11] extended the paradigm of image contrastive learning established by BYOL [20] by designing different temporal views to circumvent the challenges associated with negative sample construction in contrastive learning. Nonetheless, there exists the potential for performance degradation of the model when the patch settings are not appropriately configured.
Additionally, they either do not explicitly model inter-channel dependencies [11] or rely solely on basic MLPs or Transformers that focus on temporal information [18], mainly considering point-level features [19]. However, with the advancement of TSAD evaluation metrics, an increasing number of interval-based evaluation indicators [21, 22, 23, 24] have garnered research attention. Nevertheless, prior studies predominantly concentrated on the F1 with point ajustment [17], thereby neglecting the detection of interval anomalies. Consequently, there remains considerable potential for enhancing the detection capabilities for interval anomalies in current research [24].
@@
Furthermore, the previous TSAD algorithms based on contrastive learning [17, 11, 12, 18] lack in-depth exploration from a theoretical perspective and fail to provide more valuable intuitions for future research. In contrast, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points. This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not. This fundamental principle forms the basis of our proposed anomaly detection method. Simultaneously, to further explore this concept, we conducted both experimental and theoretical analyses to investigate the underlying principles of this approach, thereby addressing the challenges from various perspectives.
@@
In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection. Specifically, proposed Multi-scale Patching and Embedding decomposes the input data along the temporal dimension into non-overlapping patches, extracting semantically richer features beyond simple point-level characteristics. Besides, various submodels address various time scales, each focusing on different patch sizes. Patch Mixer Encoders are introduced to learn the complex relationships within and between patches and channels. A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities. At the same time, the lightweight model configuration enables PatchAD to better tackle Challenge 2, while the intricate encoder design enhances the model’s feature learning capacity to confront Challenge 1.
In detail, this work makes the following contributions:
@@
Figure 2: Model Params, FLOPs, and Latency.
(1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method. PatchAD demonstrates exceptional performance across eight benchmark datasets, exceeding 30 baselines across a range of evaluation metrics, including commonly used metrics and interval anomaly indicators.
(2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models. Moreover, we conducted an analysis of the model’s scalability from both experimental and theoretical perspectives.
(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by practical and theoretical analyses that elucidate the model’s underlying principles and design choices.
II Preliminaries
II-A Problem Formulation
@@
II-B Time Series Anomaly Detection
@@
The time series anomaly detection methods encompass statistical approaches, classical machine learning methods, and deep learning techniques [25]. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [26] and Support Vector Data Description (SVDD) [27]. However, machine learning methods often struggle to extract meaningful features and detect anomalies when confronted with high-dimensional, complex, and voluminous time series data, ultimately resulting in suboptimal performance [4, 28].
Deep learning methods include OmniAnomaly [9], USAD [19], LSTM-VAE [8], and DeepAnT [29], which are trained on normal data and detect anomalies by inference using reconstruction errors. They often rely on reconstruction assumptions, wherein the model’s ability to restore anomalous data diminishes after being trained on normal data. Previous studies [30, 31] have indicated that this prior becomes less applicable when the model’s reconstruction capability is excessively strong, leading to a decline in anomaly detection performance.
Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [32]. While [33] designed six different anomaly injection methods to train the model in a self-supervised approach, [34] utilized outlier exposure and mixup to construct a self-supervised task. However, due to the distributional discrepancies between real and synthetic data, this results in subpar performance when confronted with real-world data, as observed in [33, 34].
Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [35] introduced sequence variational autoencoder model based on smoothness constraints, whereas [17] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance. Recently approaches based on diffusion models and Large language model have also been proposed [36, 16]. Although significant advancements have been made in the performance of time series anomaly detection methods, most of these approaches are based on Transformers or sophisticated attention models [17, 11, 36, 16]. This reliance results in models that are not sufficiently lightweight and exhibit slower computational efficiency, rendering them unsuitable for certain real-world scenarios that require rapid detection.
II-C MLP Mixer & Contrastive Learning for Time Series Analysis
@@
The MLP Mixer initially proposed as a novel architecture for computer vision [37], has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like [38, 39, 40, 41] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. For example, PatchTST [39] employs a single-scale patch-based MLP as its core architecture, whereas the work in [41] applies an MLP to capture frequency-domain information in time series data for predictive modeling. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time series forecasting (LSTF). Constrained by space, further detailed descriptions about Mixer are provided in Appendix -A1.
@@
Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs [20, 42, 12]. Additionally, methods like BYOL [20] and SimSiam [42] have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as [12] and [43] have explored integrating contrastive learning into time series representation learning. TS2Vec [44] performs contrastive learning in a hierarchical way over augmentation views. [17, 11], on the other hand, combine self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection. In consideration of space limits, technical details are provided in Appendix -A2.
III Proposed Method
III-A Overall Architecture
@@
III-B Multi-Scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. Patching divides the time series window into smaller patches [39]. Multi-scale Patching enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$ [39]. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ (\mathbb{R}^{C\times T}\rightarrow\mathbb{R}^{C\times N\times P}),$ (1)
where $\operatorname{PE}(\cdot)$ is positional embedding [17]. PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. To make feature extraction and learning easier, these are used to embed the intra-patching and inter-patching information into the $D$-dimensional space, respectively. And $\operatorname{VE}(\cdot)$ is a single linear layer.
$\mathcal{N}$ $=\operatorname{VE}(\mathcal{X}),\ (\mathbb{R}^{C\times N\times P}\rightarrow\mathbb{R}^{C\times N\times D}),$ (2)
$\mathcal{P}$ $=\operatorname{VE}(\mathcal{X}),\ (\mathbb{R}^{C\times N\times P}\rightarrow\mathbb{R}^{C\times P\times D}).$
For the sake of brevity and clarity, the above describes the process of single-scale patching. In the last part of this chapter, we describe how multi-scale patching is implemented (see Sec. III-G).
III-C Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. The Multi-Layer Perceptron (MLP) consists of two fully connected (FC) layers, a normalization layer and a ReLU activation, and utilizes a residual connection from input to output. The MLP can be represented as:
$\operatorname{MLP}(\mathcal{X})=\operatorname{FC}(\operatorname{ReLU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (3)
The MLP Mixers of different types utilize MLP to model the feature variations along different dimensions, which can be represented as:
$\operatorname{Mixer}(\mathcal{X},dim=i)$ (4)
$=\operatorname{Transpose}(\operatorname{MLP}(\operatorname{Transpose}(\mathcal{X},\text{dim = i})),\text{dim = i}),$
where $\operatorname{Transpose}(\cdot,\text{dim = i})$ denotes the operation of swapping the $i$-th dimension with the last dimension. The purpose of this is to enable the MLP to effectively capture the features along different dimensions. The following are the different types of MLP Mixers utilized in PatchAD:
@@ III-C Patch Mixer Encoder
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation. Specifically, for each layer of the Patch Mixer Encoder, there exist two types of input features: inter-patch, denoted as $\mathcal{N}$, and intra-patch, represented as $\mathcal{P}$.
Next, we elaborate on the workflow of four types of mixers in detail. Initially, these features are processed by the shared-parameter Channel Mixer, expressed as follows:
[…]
Similarly, the Inter Mixer and Intra Mixer utilize the outputs of the channel mixer as their respective inputs, yielding:
[…]
Here, $\operatorname{InterMixer}$ and $\operatorname{InraMixer}$ denote the functions of the Inter Mixer and Intra Mixer, respectively. It is important to note that although both operate on feature variation at dim=2, the differing shapes of the input tensors result in $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$. Consequently, $\operatorname{InterMixer}$ processes $\mathcal{N}$ along dimension $N$ to learn detailed features of patches, whereas $\operatorname{IntraMixer}$ processes $\mathcal{P}$ along dimension $P$ to acquire global knowledge of patches. Secs. V-A and V-B provide a detailed experimental and theoretical analysis of the nature of learning in both cases.
Subsequently, to enable the model to learn low-dimensional invariant information from different types of features and thereby enhance the generalization capability of PatchAD, inspired by [45], the MixRep Mixer is employed to embed the two distinct types of features into a unified representation space, expressed as follows:
[…]
$\mathcal{N}=\operatorname{MixRepMixer}(\operatorname{InterMixer}(\operatorname{ChannelMixer}(\mathcal{N}))),$ (8)
$\mathcal{P}=\operatorname{MixRepMixer}(\operatorname{IntraMixer}(\operatorname{ChannelMixer}(\mathcal{N}))).$
In summary, the four mixer workflows correspond to Figure 3(b).
In the end, two separate simple MLPs, e.i., $\operatorname{Rec_{1}}$ and $\operatorname{Rec_{2}}$ are utilized to reconstruct the original data, which are then combined.
$\hat{\mathcal{X}}_{1}=\operatorname{Rec_{1}}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec_{2}}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (9)
Simultaneously, the structures of $\operatorname{Rec_{1}}$ and $\operatorname{Rec_{2}}$ are identical. Initially, the last two channels of $\mathcal{N}$ and $\mathcal{P}$ are flattened using $\operatorname{Flatten}(\cdot)$. Subsequently, the features are reconstructed into the original time series $\mathcal{X}$ through a MLP. This process can be represented as:
$\operatorname{Rec_{1}}(\mathcal{N})$ $=\operatorname{MLP}(\operatorname{Flatten}(\mathcal{N})),$ (10)
$\operatorname{Rec_{2}}(\mathcal{P})$ $=\operatorname{MLP}(\operatorname{Flatten}(\mathcal{P})).$
III-D Dual Project Constraint